OpenStudy (anonymous):

in triangle DEF, DE= 15 and measurement of f is 32. e is 90 degrees. find EF to the nearest tenth

OpenStudy (anonymous):

is 32 the angle at F?

OpenStudy (amistre64):

D------ 15 -----E 90 / / / / / ?? / / 32 / F

OpenStudy (amistre64):

tan(32) = 15/EF EF = 15/tan(32)

OpenStudy (amistre64):

EF = 24.005

OpenStudy (amistre64):

worst drawing ever!!

OpenStudy (anonymous):

Could also use the law of sines =]